In a series of 30 racquetball games played to date, larry has won 10 giving him a winning average of 33.3% if he continues to play what interval describes the number of games he must now win in a row to have an overall winning average greater than 50 %
Answer by Alan3354(69443)

In a series of 30 racquetball games played to date, larry has won 10 giving him a winning average of 33.3% if he continues to play what interval describes the number of games he must now win in a row to have an overall winning average greater than 50 %
=================
He has to win W games.
(10+W)/(30+W) = 1/2
W+30 = 2(W+10) = 2W+20
W = 10
================
--> 20 out of 40 = 50%